# Local stability guarantees for data-driven quadratically nonlinear models

### v3

@@ III.2 Proposed extended trapping SINDy algorithm
Similarly to Sawant et al[42] and Erichson et al[43], we can guarantee local stability and promote large stable radii for data-driven models with nonlinearities that do not preserve the energy by minimizing the destabilizing effects of the cubic terms in $\dot{v}(\bm{y})$. The Sawant et al[42] optimization is approximately,
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (38)
@@
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (39)
@@
We now additionally promote locally stable models without minimizing the nonlinear effect of the systems. Recall from Thm. 2 that under certain conditions, linear-quadratic models can be locally bounded and exhibit a trapping region even if the hard constraint of the energy-preserving structure is relaxed. For ease of notation, we introduce $Q^{\prime}_{ijk}=Q_{ijk}+Q_{jki}+Q_{kij}$, $\bm{H}_{0}=\mathcal{T}(\bm{Q}^{\prime})$ and $\tilde{\bm{H}}_{0}=\bm{P}^{-\frac{1}{2}}\mathcal{T}\bm{H}_{0}(\bm{P}^{-\frac{1}{2}}\otimes\bm{P}^{-\frac{1}{2}})$. The objective function now reads,
$\begin{gathered}\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\left[\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}+\frac{\bm{\lambda}_{1}(\tilde{\bm{A}}^{S})}{\eta}\right]\\ \text{s.t.}\quad\|\tilde{\bm{H}}_{0}\|_{F}=\epsilon_{Q},\end{gathered}$ (40)
@@
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (41)
@@
Eq. (41) is quite similar to the objective function used in Ouala et al[17]. This optimization can be understood as maximizing the stability radius about the origin at $\bm{m}$ by promoting energy-preserving structure in $\bm{Q}$, instead of maximizing the stability radius about the radius by minimizing the norm of $\bm{Q}$ as in Sawant et al[42], which makes the nonlinearities weak. Considering that there is no “true zero” due to the existence of machine precision in computer arithmetic; even the resulting models from Eq. (37) will not be truly globally bounded numerically. A description of each of the hyperparameters $\gamma,\eta$ and $\beta$ is provided in Table 1.
Notice that we still need $\tilde{\bm{A}}^{S}$ to be Hurwitz to get the estimated stability radius, and the largest eigenvalue $\lambda_{1}$ of $\tilde{\bm{A}}^{S}$ determines the stability radius directly. By additionally optimizing for the Lyapunov matrix $\bm{P}$, as in Goyal et al[18], we may be able to better control the magnitude of $\lambda_{1}$.
@@ III.3 Interpretation of the loss term
Instead of minimizing the $\|\bm{H}\|_{F}$, Eq. (41) minimizes $\|\tilde{\bm{H}}_{0}\|_{F}$. To explain further, we do not promote the local stability of the models by weakening the nonlinearities of the system, but by enforcing a weakly quadratic energy-preserving structure. Moreover, we here use this optimization to find the $\rho_{+}$ and $\rho_{-}$, which guarantee the stability of the model based on Thm. 2. Specifically, the size of the domain of boundedness $\rho_{+}$ and the size of the trapping region $\rho_{-}$ is derived from the optimized $\|\tilde{\bm{H}}_{0}\|_{max}$. While Ouala et al[17] make a number of important contributions and extensions to neural-network-based system identification with stability, they are missing the $\rho_{+}$ and $\rho_{-}$ estimates derived in the present work that quantify the stability regions.
We can also reproduce the results from the work by Kaptanoglu et al[2] through Eq. 41, which means our method could also promote global stability in quadratically nonlinear models. As what we show in Sec. II.5, global stability of the models can be achieved, theoretically, by reducing the loss of $\|\tilde{\bm{H}}_{0}\|_{F}$ to machine precision.
@@ IV Results
VallisElNino 4 0.022 $10^{-9}$ $10^{2}$ $6.5\times 10^{-9}$ $2.4\times 10^{3}$ $2.3\times 10^{7}$ $2.4\times 10^{3}$0.1 0.7 $10^{-5}$
Von Karman (enstrophy) $5$ $0.1$ $10^{-10}$ $10^{4}$ $5.9\times 10^{-10}$ $4\times 10^{3}$ $2.9\times 10^{8}$ $2.4\times 10^{3}$ $-0.12$
@@ IV.1 Noisy Lorenz attractor
Figure 5: A collection of Lorenz attractors with additive Gaussion noise. Left: Lorenz trajectories with noise for training (red) and the corresponding extended trapping SINDy models (black, solid). Right: Lorenz trajectories without noise for test (blue) and their prediction by extended trapping SINDy models (black, dashed).
@@ IV.2 Assorted locally stable models from the dysts database
The dysts dataset [45] is a collection of more than 100 chaotic systems ideal for benchmarking system identification methods [46]. Four systems with the energy-preserving structure in the quadratic terms (and effective nonlinearity) are picked for benchmark testing as shown in Fig 6. The datasets of all systems are normalized so we can compare the performance of one set of combinations of hyperparameters on different dynamical systems. The algorithm successfully recognizes the attractors and performs well on promoting stability in modeling these systems. Table 2 also gives the normalized trapping region size of each system.
The finance chaotic system shown in Fig. 6(a) is defined in Cai et al[47], and described as
@@
Figure 6: A collection of benchmark test on dysts database. Trajectories of (a) A finance chaotic system, (b) Hadley cell dynamics, (c) Stochastic Lorenz–Stenflo system, (d) El Niño phenomenon for training (yellow) and locally stable trapping SINDy models (blue) capturing the strange attractors on these training data, producing large stable radii for all four systems that satisfy the totally-symmetric quadratic coefficient constraints.
@@ IV.3 Von Karman vortex street using the enstrophy
The fluid wake behind a bluff body is characterized by a periodic vortex shedding phenomenon known as a von Kármán street. The two-dimensional incompressible flow past a cylinder is a typical example of such behavior, and has been a benchmark problem for Galerkin models for decades. Early Galerkin models of vortex shedding, based on a POD expansion about the mean flow, captured the oscillatory behavior but were structurally unstable [51]. This was later resolved by Noack et al[3], in which an 8-mode POD basis was augmented with a ninth “shift mode” parameterizing the mean flow deformation. This approach was later formalized with a perturbation analysis of the flow at the threshold of bifurcation [52].
Figure 7: Five-dimensional models, integrated from three random initial conditions (color lines), for the von Karman street from the traditional trapping SINDy using the energy as a Lyapunov function, the analytic POD-Galerkin method, and the locally stable trapping SINDy using the enstrophy. The POD modes from the direct numerical simulation data were used for training and are shown in black lines.
@@
Recent work by Loiseau et al[41, 53, 54] bypassed the Galerkin projection step by using the SINDy algorithm to directly identify the reduced-order dynamics. This approach has been shown to yield compact, accurate models for low-dimensional systems ($r=2$ or $3$), but preserving accuracy and stability for higher-dimensional systems remains challenging. Higher-dimensional regression problems often become ill-conditioned; for example, in the cylinder wake example, the higher modes 3-8 are essentially harmonics of the driving modes 1-2, and so it is difficult to distinguish between the various polynomials of these modes during regression. Because these higher harmonics are driven by modes 1-2, the 3D constrained quadratic SINDy model with modes 1-2 plus the shift mode from Loiseau et al[41] already performs well enough to capture the energy evolution with minor overshoot and correct long-time behavior.
@@
computed using the DNS, is used as the Lyapunov matrix in Eq. (24). Both trapping SINDy models show improved accuracy with respect to the POD-Galerkin model in recovering the true limit cycle dynamics in Fig. 7.
@@ IV.4 Lid-cavity flow
In fluid mechanics, the lid-driven cavity is commonly used as a benchmark for testing numerical methods [55]. A viscous fluid is in a confined geometry of a 2D rectangular container, and its motion is driven by the tangential moving of a bounding wall. The fluid motion is governed by incompressible Navier-Stokes equations. We then apply the model reduction method previously mentioned, resulting in a POD-Galerkin model. In this work, we select the first 6 POD modes to build a data-driven model.
The flow considered in the present work is the incompressible shear-driven cavity flow at $Re=7500$. The computational domain and boundary conditions are the same as in Callaham et al[56]. We also use the same data from direct numerical simulation (DNS) as in the work. Then the extended trapping SINDy algorithm is applied to compare with direct numerical simulation, 6-modes POD-Galerkin model, 16-modes POD-Galerkin model and a naive SINDy model. As shown in Fig. 8, the extended trapping SINDy model outperforms all other presented data-driven models and indicates strong stability compared to all others, even though the $\tilde{\bm{A}}^{S}$ from the extended trapping SINDy is not fully promoted to negative-definite. The $\tilde{\bm{A}}^{S}$ has one positive eigenvalue which is similar as the one in the $50\%$ noisy Lorenz example in Sec. IV.1. Further hyperparameter tuning might facilitate a fully provably stable model but we found empirically this model already works well for the lid cavity flow.
Figure 8: Lid-driven cavity. $K$ denotes the total kinetic energy. $t$ is the normalized time. Extended trapping SINDy and naive SINDy both use the first 6 POD modes. The naive SINDy model goes unstable immediately and POD-Galerkin models go unstable gradually, while the locally stable model captures the stable dynamics using the same or fewer number of modes.
@@ V Conclusion
In Sec. III, an extended trapping SINDy algorithm was proposed to promote local stability of data-driven quadratically nonlinear models. One of the guarantees on performance needed to use this algorithm is given by the local stability estimates from the theorem proposed in Sec. II.4. The algorithm performs well on tasks such as system identification and simulation with ROMs without requiring the nonlinearities of the dynamical system to be small to have these guarantees. Successfully promoting the local stability of data-driven models requires a combination of well designed algorithms, methods of analysis with physical insight and interpretations, and local or global stability guarantees. The work of this paper was motivated, in part, by the idea that tools that extract models from data should guarantee the stability of models that capture the dominant behavior of bounded physical systems like fluids which can be described by Navier-Stokes equations. We observe many situations where the characteristics of the dynamics or stability/instability conditions are unclear in our presented example results and unpublished experiment, which includes that there exists one positive eigenvalue of $\tilde{\bm{A}}^{S}$ and that no stability guarantee is given by the Thm. 2 in some models but yet to find any unstable solution in neither of these situations. This leads to a number of possible directions for future work that would extend the research presented in this paper.
We train our extended trapping SINDy algorithm on a single trajectory for all of the examples presented in this work. And for most data-driven methods, the performance can be improved by training larger datasets, meaning processing data from multiple trajectories.
As mentioned above in Sec. IV.1 and Sec. IV.4, although the matrix $\tilde{\bm{A}}^{S}$ from Eq. (16b) is not strictly Hurwitz, we observe robust local stability of these SINDy models. The initial conditions that leads to an unstable solution are still not clear in this work. Even though we have given sufficient conditions for attracting trapping regions to exist in quadratically nonlinear systems and conservative estimates about the size of the local stability domain, an unbounded trajectory has not been yet found when the Hurwitz property is slightly broken. The inability to find a $\bm{m}$ such that $\tilde{\bm{A}}^{S}$ is Hurwitz can always be an issue of a suboptimal minima in the optimization problem, but dynamical explanations may be required too.
@@
It roughly follows from this reasoning that if we promote models such that more and more eigenvalues of $\tilde{\bm{A}}^{S}$ become negative, this will contribute to the stability whether or not $\tilde{\bm{A}}^{S}$ is able to become fully Hurwitz. This is precisely our empirical finding in both the local and global versions of the trapping SINDy algorithm.
Future numerical work includes improving the strategy of enforcing a Hurwitz matrix $\tilde{\bm{A}}^{S}$ during optimization. The current strategy to enforce a Hurwitz matrix $\tilde{\bm{A}}^{S}$ is to solve $\bm{\xi}$ and $\bm{m}$ separately in each iteration so that each process is convex. However, the loss term for making a Hurwitz $\tilde{\bm{A}}^{S}$ inevitably makes the loss function nonconvex. The convergence properties are unclear as well. A clever strategy for enforcing the Hurwitz property was adopted recently in Goyal et al[18]. Alternatively, with the help of modern powerful computational platforms, the nonconvex problems can be better solved by implementing this method in deep learning. So developing a deep neural network for this method as in Ouala et al[17] is another promising line of future work.
Lastly, Kramer raises an optimization-based approach to enlarge the estimate of the stability region for a given Lyapunov matrix $\bm{P}$ [19]. From Thm. 2 and Lyapunov function (24), we can observe that it is possible to control the magnitudes of the eigenvalues of $\tilde{\bm{A}}^{S}$ using different Lyapunov matrices $\bm{P}$. This implies that some $\bm{P}$ can enlarge the the guaranteed stability region by forming $\tilde{\bm{A}}^{S}$ with an even greater largest eigenvalue so that $\rho_{+}$ is larger. Searching for an optimal $\bm{P}$ through data-driven methods as in Goyal et al[18] and machine learning algorithms is promising to extend the algorithm in this work even further.

### v4

@@ III.2 Proposed extended trapping SINDy algorithm
Similarly to Sawant et al. [42] and Erichson et al. [43], we can guarantee local stability and promote large stable radii for data-driven models with nonlinearities that do not preserve the energy by minimizing the destabilizing effects of the cubic terms in $\dot{v}(\bm{y})$. The Sawant et al. [42] optimization is approximately,
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}^{\prime}}\biggr[$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (38)
@@
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}^{\prime}}\biggr[$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (39)
@@
We now additionally promote locally stable models without minimizing the nonlinear effect of the systems. Thm. 2 enables local stability guarantees for quadratic systems with weakly energy-preserving nonlinearities. Recall that under certain conditions, linear-quadratic models can be locally bounded and exhibit a trapping region even if the hard constraint of the energy-preserving structure is relaxed. For ease of notation, we introduce $Q^{\prime}_{ijk}=Q_{ijk}+Q_{jki}+Q_{kij}$, $\bm{H}_{0}=\mathcal{T}(\bm{Q}^{\prime})$ and $\tilde{\bm{H}}_{0}=\bm{P}^{-\frac{1}{2}}\mathcal{T}\bm{H}_{0}(\bm{P}^{-\frac{1}{2}}\otimes\bm{P}^{-\frac{1}{2}})$. The objective function now reads,
$\begin{gathered}\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}^{\prime}}\left[\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}+\frac{\bm{\lambda}_{1}(\tilde{\bm{A}}^{S})}{\eta}\right]\\ \text{s.t.}\quad\|\tilde{\bm{H}}_{0}\|_{F}=\epsilon_{Q},\end{gathered}$ (40)
@@
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}^{\prime}}\biggr[$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (41)
@@
Eq. (41) is quite similar to the objective function used in Ouala et al. [17]. This optimization can be understood as maximizing the stability radius about the origin at $\bm{m}$ by promoting energy-preserving structure in $\bm{Q}$, instead of maximizing the stability radius about the radius by minimizing the norm of $\bm{Q}$ as in Sawant et al. [42], which makes the nonlinearities weak. When $\|\tilde{\bm{H}}_{0}\|_{F}^{2}$ is reduced to 0, global stability is obtained and strict energy-preserving constraints can be found in $\bm{Q}$. In other words, the constrained optimization in [2] is achieved using this unconstrained optimization. Moreover, by introducing the Lyapunov matrix $\bm{P}$, we generalize the energy-preserving structure which is restricted to the kinetic energy of the flow in the very original trapping SINDy algorithm. Considering that there is no “true zero” due to the existence of machine precision in computer arithmetic; even the resulting models from Eq. (37) will not be truly globally bounded numerically. A description of each of the hyperparameters $\gamma,\eta$ and $\beta$ is provided in Table 1.
Notice that we still need $\tilde{\bm{A}}^{S}$ to be Hurwitz to get the estimated stability radius, and the largest eigenvalue $\lambda_{1}$ of $\tilde{\bm{A}}^{S}$ determines the stability radius directly. By additionally optimizing for the Lyapunov matrix $\bm{P}$, as in Goyal et al. [18], we may be able to better control the magnitude of $\lambda_{1}$.
@@ III.3 Interpretation of the loss term
Instead of minimizing the $\|\bm{H}\|_{F}$, Eq. (41) minimizes $\|\tilde{\bm{H}}_{0}\|_{F}$. To explain further, we do not promote the local stability of the models by weakening the nonlinearities of the system, but by enforcing a weakly quadratic energy-preserving structure. Moreover, we here use this optimization to find the $\rho_{+}$ and $\rho_{-}$, which guarantee the stability of the model based on Thm. 2. Specifically, the size of the domain of boundedness $\rho_{+}$ and the size of the trapping region $\rho_{-}$ is derived from the optimized $\|\tilde{\bm{H}}_{0}\|_{max}$. While Ouala et al. [17] make a number of important contributions and extensions to neural-network-based system identification with stability, they are missing the $\rho_{+}$ and $\rho_{-}$ estimates derived in the present work that quantify the stability regions.
We can also reproduce the results from the work by Kaptanoglu et al. [2] through Eq. 41, which means our method could also promote global stability in quadratically nonlinear models. As what we show in Sec. II.5, global stability of the models can be achieved, theoretically, by reducing the loss of $\|\tilde{\bm{H}}_{0}\|_{F}$ to machine precision.
@@ IV Results
VallisElNino 4 0.022 $10^{-9}$ $10^{2}$ $6.5\times 10^{-9}$ $2.4\times 10^{3}$ $2.3\times 10^{7}$ $2.4\times 10^{3}$ -0.1 0.7 $10^{-5}$
Von Karman (enstrophy) 55 $0.1$ $10^{-10}$ $10^{4}$ $5.9\times 10^{-10}$ $4\times 10^{3}$ $2.9\times 10^{8}$ $2.4\times 10^{3}$ 0.12-0.12
@@ IV.1 Noisy Lorenz attractor
Figure 5: A collection of Lorenz attractors with additive Gaussion noise. Left: Lorenz trajectories with noise for training (black), the corresponding extended trapping SINDy models (red) and the corresponding base SINDy models (green). Right: Lorenz trajectories without noise for test (black, their prediction by extended trapping SINDy models (red) and prediction by base SINDy models (green).
@@ IV.2 Assorted locally stable models from the dysts database
The dysts dataset [45] is a collection of more than 100 chaotic systems ideal for benchmarking system identification methods [46]. Four systems with the energy-preserving structure in the quadratic terms (and effective nonlinearity) are picked for benchmark testing as shown in Fig 6. While these systems theoretically possess lossless nonlinearities, they cannot be maintained numerically. Nevertheless, the dataset remains valid for testing the extended algorithm and deriving stability estimates for the resulting numerical models. The datasets of all systems are normalized so we can compare the performance of one set of combinations of hyperparameters on different dynamical systems. The algorithm successfully recognizes the attractors and performs well on promoting stability in modeling these systems as well as the base SINDy, which is because we use clean data to train each model and the systems are all well-defined with large stability domain. However, base SINDy usually does not perform well on noisy data as we have shown in the previous example. Table 2 also gives the normalized trapping region size of each system.
The finance chaotic system shown in Fig. 6(a) is defined in Cai et al. [47], and described as
@@
Figure 6: A collection of benchmark test on dysts database. Trajectories of (a) a finance chaotic system, (b) Hadley cell dynamics, (c) Stochastic Lorenz–Stenflo system, (d) El Niño phenomenon for testing (black). Locally stable trapping SINDy models (red) capture and predict the strange attractors on these testing data, producing large stable radii for all four systems that satisfy the totally-symmetric quadratic coefficient constraints, each with a comparing prediction from accordingly base SINDy model (green).
@@ IV.3 Von Karman vortex street using the enstrophy
The fluid wake behind a bluff body is characterized by a periodic vortex shedding phenomenon known as a von Kármán street. The two-dimensional incompressible flow past a cylinder is a typical example of such behavior, and has been a benchmark problem for Galerkin models for decades. Early Galerkin models of vortex shedding, based on a POD expansion about the mean flow, captured the oscillatory behavior but were structurally unstable [51]. This was later resolved by Noack et al. [3], in which an 8-mode POD basis was augmented with a ninth “shift mode” parameterizing the mean flow deformation. This approach was later formalized with a perturbation analysis of the flow at the threshold of bifurcation [52].
Figure 7: Five-dimensional models, integrated from three random initial conditions (color lines), for the von Karman street from the traditional trapping SINDy using the energy as a Lyapunov function, the analytic POD-Galerkin method, and the locally stable trapping SINDy using the enstrophy. The POD modes from the direct numerical simulation data were used for training and are shown in black lines. Each figure is plotted over the same axis dimensions and each row of the subplots shares the same combination of coordinates/modes.
@@
Recent work by Loiseau et al. [41, 53, 54] bypassed the Galerkin projection step by using the SINDy algorithm to directly identify the reduced-order dynamics. This approach has been shown to yield compact, accurate models for low-dimensional systems ($r=2$ or $3$), but preserving accuracy and stability for higher-dimensional systems remains challenging. Higher-dimensional regression problems often become ill-conditioned; for instance, in the cylinder wake example, the higher modes 3-8 are essentially harmonics of the driving modes 1-2, and so it is difficult to distinguish between the various polynomials of these modes during regression. Because these higher harmonics are driven by modes 1-2, the 3D constrained quadratic SINDy model with modes 1-2 plus the shift mode from Loiseau et al. [41] already performs well enough to capture the energy evolution with minor overshoot and correct long-time behavior.
@@
computed using the direct numerical simulation (DNS), is used as the Lyapunov matrix in Eq. (24). Both trapping SINDy models show improved accuracy with respect to the POD-Galerkin model in recovering the true limit cycle dynamics in Fig. 7.
@@ IV.4 Lid-cavity flow
Figure 8: Lid-driven cavity, where $K$ denotes the total kinetic energy, and $t$ is the normalized time. Extended trapping SINDy and naive SINDy both use the first 6 POD modes. The lower subplots is a zoom-in of the upper subplots between time 30 to 60.
In fluid mechanics, the lid-driven cavity is commonly used as a benchmark for testing numerical methods [55]. A viscous fluid is in a confined geometry of a 2D rectangular container, and its motion is driven by the tangential moving of a bounding wall. The fluid motion is governed by incompressible Navier-Stokes equations. We then apply the model reduction method previously mentioned, resulting in a POD-Galerkin model. In this work, we select the first 6 POD modes to build a data-driven model. The flow considered in the present work is the incompressible shear-driven cavity flow at $Re=7500$. The computational domain and boundary conditions are the same as in Callaham et al. [56]. We also use the same data from DNS as in the work. Then the extended trapping SINDy algorithm is applied to compare with direct numerical simulation, 6-modes POD-Galerkin model, 16-modes POD-Galerkin model and a naive SINDy model. Since this system lacks lossless nonlinearities, applying the original trapping SINDy method for comparison would be inappropriate. As shown in Fig. 8, the extended trapping SINDy model outperforms all other presented data-driven models and indicates strong stability compared to all others, even though the $\tilde{\bm{A}}^{S}$ from the extended trapping SINDy is not fully promoted to negative-definite. The naive SINDy model goes unstable immediately and POD-Galerkin models go unstable gradually, while the locally stable model captures the stable dynamics using the same or fewer number of modes. One may also notice the increasing phase shift between the DNS results and the trapping SINDy model, along with the reduction in total kinetic energy predicted by the trapping SINDy model. Both observations can likely be attributed to the same cause, which is that the trapping model takes only the first 6 POD modes to simulate the dynamics, and this truncation inevitably leads to a slight loss of energy. Additional results are provided in Appendix B.2. The $\tilde{\bm{A}}^{S}$ has one positive eigenvalue which is similar as the one in the $50\%$ noisy Lorenz example in Sec. IV.1. Further hyperparameter tuning might facilitate a fully provably stable model but we found empirically this model already works well for the lid cavity flow.
@@ V Conclusion
In Sec. III, an extended trapping SINDy algorithm was proposed to promote local stability of data-driven quadratically nonlinear models. One of the guarantees on performance needed to use this algorithm is given by the local stability estimates from the theorem proposed in Sec. II.4. The algorithm performs well on tasks such as system identification and simulation with ROMs without requiring the nonlinearities of the dynamical system to be small to have these guarantees. Successfully promoting the local stability of data-driven models requires a combination of well designed algorithms, methods of analysis with physical insight and interpretations, and local or global stability guarantees. The work of this paper was motivated, in part, by the idea that tools that extract models from data should guarantee the stability of models that capture the dominant behavior of bounded physical systems like fluids which can be described by Navier-Stokes equations. We observe many situations where the characteristics of the dynamics or stability/instability conditions are unclear in our presented example results and unpublished experiment, which includes that there exists one positive eigenvalue of $\tilde{\bm{A}}^{S}$ and that no stability guarantee is given by the Thm. 2 in some models but yet to find any unstable solution in neither of these situations.
@@
It roughly follows from this reasoning that if we promote models such that more and more eigenvalues of $\tilde{\bm{A}}^{S}$ become negative, this will contribute to the stability whether or not $\tilde{\bm{A}}^{S}$ is able to become fully Hurwitz. This is precisely our empirical finding in both the local and global versions of the trapping SINDy algorithm. This leads to a number of possible directions for future work that would extend the research presented in this paper.
One direction in the future is to find the initial conditions that leads to an unstable solution that are still not clear in this work. Even though we have given sufficient conditions for attracting trapping regions to exist in quadratically nonlinear systems and conservative estimates about the size of the local stability domain, an unbounded trajectory has not been yet found when the Hurwitz property is slightly broken. The inability to find a $\bm{m}$ such that $\tilde{\bm{A}}^{S}$ is Hurwitz can always be an issue of a suboptimal minima in the optimization problem, but dynamical explanations may be required too.
Future numerical work includes improving the strategy of enforcing a Hurwitz matrix $\tilde{\bm{A}}^{S}$ during optimization. The current strategy to enforce a Hurwitz matrix $\tilde{\bm{A}}^{S}$ is to solve $\bm{\xi}$ and $\bm{m}$ separately in each iteration so that each process is convex. However, the loss term for making a Hurwitz $\tilde{\bm{A}}^{S}$ inevitably makes the loss function nonconvex. The convergence properties are unclear as well. A clever strategy for enforcing the Hurwitz property was adopted recently in Goyal et al. [18]. Alternatively, with the help of modern powerful computational platforms, the nonconvex problems can be better solved by implementing this method in deep learning. So developing a deep neural network for this method as in Ouala et al. [17] is another promising line of future work.
Lastly, Kramer raises an optimization-based approach to enlarge the estimate of the stability region for a given Lyapunov matrix $\bm{P}$ [19]. However, optimizing the Lyapunov matrix $\bm{P}$ could expand the stability radius $\rho_{+}$ as well. Thm. 2 suggests that tuning $\bm{P}$ to reshape $\tilde{\bm{A}}^{S}$’s eigenvalues may achieve this, as demonstrated in Sec. III.2. Searching for an optimal $\bm{P}$ through data-driven methods as in Goyal et al. [18] and machine learning algorithms is promising to extend the algorithm in this work even further.